You want a seat on the board of directors of Zeph, Inc. The company has 315,000 shares of stock outstanding and the stock sells for $46 per share. There are currently 4 seats up for election. If the company uses cumulative voting, how much will it cost you to guarantee that you will be elected to the board?
Answer #1

Hi

In cumulative voting even minority share holders will have their say in electing the directors.

Here Total number of shares S = 315,0000

Total numbers of directors needed N = 1

Total number of seats for election D =4

So total numbers of shares needed to control seat = S*N/(D+1) +1

= 315000*1/(4+1) +1

= 63000+1 = 63,001

Per share price = $46

So total cost to guarantee the seat = 63001*46 = $2,898,046

Thanks